a painter needs to have 24 gallons of paint that contains 30% solvent. how many gallons of paint containing 20% solvent ahd how many gallons containing 60% solvent should be mixed to obtain the desired paint?
Answer by mananth(16946)

paint containing 20% solvent :let this be x gallons
paint containing 60% solvent :this will be 24-x gallons
Mixture : 24 gallons of paint that contains 30% solvent.
...
0.2x+0.6(24-x) = 24*0.3
0.2x+14.4-0.6x=7.2
-0.4x=-7.2
divide by -0.4
x= -7.2/0.4
x=18
18 gallons of 20%
6 gallons of 60% will be required.
...
m.ananth@hotmail.ca